A shipment of five computers contains two

that slightly defective. If a retailer receives three of these computers at

random, list the elements of the sample space S using the letters D and N for

defective and non-defective computer, respectively. To each sample point assign a value x of the

random variable X representing the number of computers purchased by the

retailer which are slightly defective. FIND THE MEAN, VARIANCE AND STANDARD DEVIATION


1
Expert's answer
2022-02-08T16:32:06-0500

"P(X=0)=\frac{C_3^0*C_2^2}{C_5^3}=\frac{1}{10}."

"P(X=1)=\frac{C_3^1*C_2^2}{C_5^3}=\frac{3}{10}."

"P(X=2)=\frac{C_3^2*C_2^1}{C_5^3}=\frac{3}{5}."

"P(X=3)=0."

"\mu=\frac{1}{10}*0+\frac{3}{10}*1+\frac{3}{5}*2=\frac{3}{2}=1.5."

"\sigma^2=\frac{1}{10}*0^2+\frac{3}{10}*1^2+\frac{3}{5}*2^2-(\frac{3}{2})^2=\frac{9}{20}=0.45."

"\sigma=\sqrt{0.45}\approx 0.67."